OpenStudy (anonymous):

What is the volume of a cylinder inside a cube with side length 10 cm?

OpenStudy (anonymous):

Hi what don't you understand about this question?

OpenStudy (anonymous):

how to find the volume of the cylnder

OpenStudy (anonymous):

The volume of a cylinder = area of base x its height

OpenStudy (anonymous):

I do not think this is such an easy question. Do you know the volume of any cylinder? (height=h, radius of base=r)

OpenStudy (anonymous):

Since the base of cylinders is always a circle, you use A=(pi)r^2 to find the base area.

OpenStudy (anonymous):

This question is obviously badly worded, since the cylinder can be infinitely small sitting inside the cube. But assuming that it is telling us to find the maximum volume of the cylinder, we first need to determine the radius of the cylinder's base. Do you know how to do that?

OpenStudy (anonymous):

no i dont have any measurments for the cylinder just the cube

OpenStudy (jidnyesha):

since the cylinder is inside the cube of 10cm, the diameter of the circular crossection will be 10cm and so will be the height h=10cm and d=10cm so r= 5cm volume of cylinder=(pi)*r^2*h therefore, vol=(pi)*5^2*10 =(pi)*25*10 =(pi)*250

OpenStudy (anonymous):

Yea. It's asking you to use the measurement of the cube to determine the radius